An initial volume of 3.7 L of a gas are maintained at constant pressure while being cooled from 95 ºC to 22 ºC. What is the new volume of the gas?
1
Expert's answer
2020-12-01T09:09:56-0500

According to the law of volumes:

"\frac{V_1} {T_1} = \frac{V_2} {T_2}" for the same amount of a gas at constant pressure.


"V_{new} = V_{orginal} * \frac{T(22\ ^0C)}{T(95\ ^0C)};"


"T(t) = (273.15 + t) \ K \ ;"


"T(22\ ^0C) = (273.15 + 22) \ K = 295.15\ K;"

"T(95\ ^0C) = (273.15 + 95) \ K = 368.15\ K;"


"V_{new} = 3.7\ L * \frac{295.15\ K}{368.15\ K}\approx 2.966\ L;"


Answer: "V_{new} = 2.966\ L"